### v1

# Ferromagnetism Mechanism in a Geometrically Frustrated Triangular Lattice

Abstract
We study the emergent itinerant ferromagnetism and propose its underlying mechanism in the geometrically frustrated triangular lattice. Based on large-scale density matrix renormalization group simulations and unrestricted Hartree-Fock mean-field analysis, we identify itinerant ferromagnetic phases in the intermediate-$U$ Hubbard model with finite doping and reveal the kinetic mechanisms assisted by geometric frustration. Notably, we find that the doublon-singlon exchange process among other microscopic charge hoppings solely drives the fully polarized ferromagnetism for geometrically frustrated triangular lattice. Additionally, we establish the whole magnetic phase diagram and illustrate itinerant ferromagnetism within a finite range of electron doping for finite on-site Coulomb repulsion. The comparison of local spin correlations with recent cold-atom experiments is also discussed. Our work enhances the understanding of ferromagnetism mechanisms at intermediate coupling strength and finite doping concentrations.
Introduction.—It has been a long-standing issue in physics to understand the ferromagnetism mechanism in strongly correlated systems. As the fundamental theoretical model for diverse physics of these systems, the Hubbard model Tasaki (1998a, b); Vollhardt et al. (1999); Arovas et al. (2022); Qin et al. (2022) is initially formulated to understand the origin of ferromagnetism Gutzwiller (1963); Hubbard and Flowers (1963). The Hubbard Hamiltonian only includes the charge kinetic-energy term with the nearest-neighbor (NN) hopping amplitude $t$ and on-site Coulomb repulsion $U$,
$H=-t\sum_{\langle\mathbf{i}\mathbf{j}\rangle,\sigma}\left(c_{\mathbf{i},\sigma}^{\dagger}c_{\mathbf{j},\sigma}+\text{H.c.}\right)+U\sum_{\mathbf{i}}n_{\mathbf{i},\uparrow}n_{\mathbf{i},\downarrow}.$ (1)
Given the lack of explicit or effective magnetic interactions favoring ferromagnetic order, ferromagnetism can only arise from the delicate interplay between kinetic energy and on-site repulsion. To date, rigorous theoretical results on the ferromagnetism mechanism in the Hubbard model have only been achieved in specific limits Tasaki (1998a, b); Vollhardt et al. (1999). The Stoner criterion Stoner (1938) states that ferromagnetism occurs when the product of the density of states at the Fermi energy $D_{\mathrm{F}}$ and on-site Coulomb repulsion $U$ satisfies the condition $UD_{\mathrm{F}}>1$. However, this criterion is derived from the Hartree-Fock approximation and is known to often overpredict ferromagnetic tendencies Vollhardt et al. (1996, 1999). Two other rigorous outcomes are Nagaoka ferromagnetism Thouless (1965); Nagaoka (1966), arising from kinetic energy minimization of a single doped hole at $U=\infty$, and flat-band ferromagnetism Mielke (1991a); Tasaki (1992) with dispersionless lowest energy bands. Over the past several decades, it has attracted substantial efforts to explore ferromagnetism beyond those well-known limits Riera and Young (1989); Basile and Elser (1990); Emery et al. (1990); Shastry et al. (1990); Mielke (1991b, 1992); Hanisch et al. (1995); Wurth et al. (1996); Brunner and Muramatsu (1998); Ulmke (1998); Becca and Sorella (2001); Merino et al. (2006); Chang et al. (2010, 2011); Carleo et al. (2011); Liu et al. (2012); Schlömer et al. (2023); Davydova et al. (2023); Samajdar and Bhatt (2023); He et al. (2023); Morera and Demler (2024), particularly for the square lattices Riera and Young (1989); Basile and Elser (1990); Emery et al. (1990); Shastry et al. (1990); Wurth et al. (1996); Brunner and Muramatsu (1998); Carleo et al. (2011); Liu et al. (2012). However, revealing the ferromagnetism mechanism in the ground state of geometrically frustrated triangular lattices at intermediate $U$ and finite doping remains challenging Merino et al. (2006); Samajdar and Bhatt (2023); He et al. (2023); Morera and Demler (2024).
More recently, the realization of cold-atom quantum simulators of the triangular-lattice Hubbard model in optical lattices Yang et al. (2021); Xu et al. (2023); Lebrat et al. (2024); Prichard et al. (2024) has significantly advanced this issue. With unprecedented control over doping and coupling strength Chin et al. (2010); Lewenstein et al. (2012), the ferromagnetic correlations are directly observed Xu et al. (2023); Lebrat et al. (2024); Prichard et al. (2024) at intermediate on-site Coulomb repulsion and finite particle doping. In a different context, the transition metal dichalcogenides (TMD) moiré materials provide another avenue for simulating the triangular-lattice Hubbard model, wherein ferromagnetism has also been observed Tang et al. (2020); Ciorciaro et al. (2023); Anderson et al. (2023); Seifert and Balents (2024). These experimental advancements necessitate theoretical study of the observed ferromagnetism and its microscopic mechanisms in the geometrically frustrated triangular lattice. Unlike the intensively studied hole-doped case, experimentally observed ferromagnetism occurs at the electron-doped side, the physics of which could be quite different due to the particle-hole asymmetry Zhu and Chen (2023); Chen et al. (2022); Xu et al. (2023); Prichard et al. (2024). In the limit of nearly half filling $n$$\to$1111 ($n$ denotes electron density per site) and full filling $n$$\to$2222, the Nagaoka-type Hanisch et al. (1995); Chen et al. (2022); Morera et al. (2023); Davydova et al. (2023); Schlömer et al. (2023); Lee et al. (2023); Morera and Demler (2024) and Müller-Hartmann-type ferromagnetism Li et al. (2023) have been proposed, respectively. However, the ferromagnetism mechanism for intermediate density around $n=3/2$ is still an open issue. At $n=3/2$, the Fermi surface nesting and van Hove singularities exist at $U=0$, weak interactions would trigger instabilities towards magnetic orders with nesting wave vectors Martin and Batista (2008); Akagi and Motome (2010). However, when the weak-coupling theory fails at the strong-coupling regime Pasrija and Kumar (2016), it is still a much-needed endeavor to identify the magnetism and its mechanism.
Motivated by the above, we explore the ferromagnetism and its underlying mechanism in the triangular-lattice Hubbard model from intermediate to infinite interactions near electron doping $\delta=1/2$ (i.e., $n=3/2$, where the doping concentration is defined by $\delta\equiv n-1$). Using large-scale density matrix renormalization group (DMRG) and unrestricted Hartree-Fock (UHF) methods, we first present the emergence of ferromagnetic ground state with metallic charge behavior at $n=3/2$ as the on-site Coulomb repulsion $U$ increases from intermediate to strong and ultimately infinite values. By examining distinct microscopic hopping processes inherent in the Hubbard model at intermediate $U$, we reveal that the doublon-singlon exchange, assisted by the geometric frustration inherited from the lattice, solely contributes to the fully polarized ferromagnetism. We further confirm our conclusions by selectively deactivating each hopping process and verifying their individual impacts on itinerant ferromagnetism. Additionally, we demonstrate that itinerant ferromagnetism around $\delta=1/2$ persists within a specific range of electron doping and find the local spin correlations at zero temperature resemble those observed in a recent cold-atom experiment Xu et al. (2023).
Fig. 1: Evolution of the magnetic and electronic structure across interaction $U/t$ at electron doping $\delta=1/2$. (a) Ferromagnetic squared order parameter $m_{s}^{2}(\mathbf{\Gamma})$ as a function of $U/t$. (b) Polarization $P$ from unrestricted Hartree-Fock method with periodic boundary conditions. (c) Line-cut plot of the static spin structure factor $S(\mathbf{q})$ along momentum path through $\mathbf{\Gamma}$ point (white dashed line in inset). Labels are for different interactions $U/t$. (d) Single-particle propagator $|C(r)|$ for different $U/t$, compared to $\sim r^{-2}$ and $\sim r^{-1}$. (e) Line-cut plot of electron momentum distribution $n(\mathbf{k})$ along momentum path through the momentum $\mathbf{\Gamma}$ (blue dashed line in inset). Insets: contour plots of $S(\mathbf{q})$ in (c) and $n(\mathbf{k})$ in (e) with $U=24$, $N=24\times 6$. White hexagon: the first Brillouin zone.
Fig. 2: Microscopic hopping processes in the Hubbard model and their kinetic energy contributions for intermediate $U/t$. (a)-(c) Schematic illustration of nearest-neighbor hopping processes between (a) a doubly occupied site and an empty site, or two singly occupied sites; (b) a singly occupied site and an empty site; (c) a doubly occupied site and a singly occupied site. These processes are described in (3). (d) Schematic illustration of ring-exchange coupling. (e),(f) The kinetic energies (in blue) of different microscopic hopping processes as functions of $U/t$ for $L_{y}=4$ in (e) and $L_{y}=6$ in (f). Ferromagnetic squared order parameter $m_{s}^{2}(\mathbf{\Gamma})$ (in orange) is plotted for comparison.
Itinerant Ferromagnetism at Electron Doping $\delta=1/2$.— We first identify the nature of the ground state in the triangular-lattice Hubbard model at electron doping $\delta=1/2$ using DMRG and UHF (details see supplementary SM ).
To examine the magnetic properties of the ground state, we use DMRG to compute the static spin structure factor $S(\mathbf{q})$, defined by $S(\mathbf{q})\equiv\sum_{\mathbf{i},\mathbf{j}}\langle{\mathbf{S}_{\mathbf{i}}\cdot\mathbf{S}_{\mathbf{j}}}\rangle e^{i\mathbf{q}\cdot(\mathbf{i}-\mathbf{j})}/N$. Magnetic order is signaled by peaks at certain wave vectors $\mathbf{q}=\mathbf{q}_{0}$. The DMRG simulations are performed on the cylinder with size $N=L_{x}\times L_{y}$, where $L_{x}$ ($L_{y}$) represents the length of the cylinder (circumference). We mainly study $L_{y}=4$, $6$ and push bond dimension $D$ up to $D\approx 36,000$. Notably, large bond dimensions are necessary for distinguishing ferromagnetic and incommensurate spin-density-wave (SDW) phases (see supplementary SM ). We first focus on the evolution of the momenta $\mathbf{q}_{0}$ as a function of interaction strength $U/t$. Our DMRG results reveal that, with increasing $U/t$ , $\mathbf{q}_{0}$ progressively shifts from $\mathbf{M}$ to $\mathbf{\Gamma}$ points SM and finally fixed at the $\mathbf{\Gamma}$ beyond a critical strength $U_{c}/t$. We therefore keep track of the squared order parameter $m_{s}^{2}(\mathbf{\Gamma})\equiv S(\mathbf{\Gamma})/N$ as a function of $U/t$. As shown in Figs. 1(a), (c), when further increasing $U/t$, $m_{s}^{2}(\mathbf{\Gamma})$ gradually increases until plateauing, signaling the establishment of a stabilized ferromagnetic phase at larger $U/t$ and the evolution from partially to fully polarized phases. The saturated value of $m_{s}^{2}(\mathbf{\Gamma})$ matches the theoretically expected $\delta_{\mathbf{q},\mathbf{\Gamma}}/16$ at $n=3/2$.
To confirm the robustness of ferromagnetism in the thermodynamic limit, we further employ the unrestricted Hartree-Fock (UHF) mean-field method SM . In the UHF method, under an uncorrelated state ansatz, both the on-site densities $\langle n_{\mathbf{i},\sigma}\rangle$ and the spin-flips $\langle S_{\mathbf{i}}^{-}\rangle$ and $\langle S_{\mathbf{i}}^{+}\rangle$ are involved to factorize the Hubbard interaction
$n_{\mathbf{i},\uparrow}n_{\mathbf{i},\downarrow}\rightarrow$ $\langle n_{\mathbf{i},\uparrow}\rangle n_{\mathbf{i},\downarrow}+\langle n_{\mathbf{i},\downarrow}\rangle n_{\mathbf{i},\uparrow}-\langle n_{\mathbf{i},\uparrow}\rangle\langle n_{\mathbf{i},\downarrow}\rangle$
$-\langle S_{\mathbf{i}}^{+}\rangle S_{\mathbf{i}}^{-}-\langle S_{\mathbf{i}}^{-}\rangle S_{\mathbf{i}}^{+}+\langle S_{\mathbf{i}}^{-}\rangle\langle S_{\mathbf{i}}^{+}\rangle.$ (2)
Here $S_{\mathbf{i}}^{+}=c_{\mathbf{i},\uparrow}^{\dagger}c_{\mathbf{i}\downarrow}$ and $S_{\mathbf{i}}^{-}=c_{\mathbf{i},\downarrow}^{\dagger}c_{\mathbf{i},\uparrow}$ are spin ladder operators. We have a total of $4N$ variational parameters, obtained through iterative diagonalization of the Hartree-Fock Hamiltonian. To ensure stable convergence, we utilize direct inversion in the iterative subspace method Pulay (1980), and the convergence criterion for the energy is set to $10^{-8}t$. As shown in Fig. 1(b), the polarization $P\equiv 2Nm_{s}(\mathbf{\Gamma})$ increases gradually as $U/t$ becomes larger, and finally a full polarization state occurs at $U\simeq 4.5t$ independently of system size. This aligns with Fig. 1(a) and is also consistent with the recent Hartree-Fock analysis He et al. (2023), indicating robust fully-polarized ferromagnetism in the strong coupling regime.
Having established the ferromagnetism in the spin channel, we further examine the electronic properties in the charge channel. We compute the single-particle propagator $C(r)\equiv\sum_{\sigma}\langle c_{\mathbf{i}_{0},\sigma}^{\dagger}c_{{\mathbf{i}_{0}+r\bm{e}_{x}},\sigma}\rangle$, and the momentum distribution of the electrons $n(\mathbf{k})\equiv\sum_{\mathbf{i},\mathbf{j},\sigma}\langle c_{\mathbf{i},\sigma}^{\dagger}c_{\mathbf{j},\sigma}\rangle e^{i\mathbf{k}\cdot(\mathbf{i}-\mathbf{j})}/N$. Figure 1(d) illustrates the power-law decay of $|C(r)|$ with an exponent approximately equal to 1, identifying the gapless nature of the electrons in the ferromagnetic phase. In the inset of Fig. 1(e), the contour plot of $n(\mathbf{k})$ exhibits an abrupt change from 2 to 1 at certain momenta, as more clearly illustrated by the line-cut plot of $n(\mathbf{k})$ through the momentum $\mathbf{\Gamma}$ in the main panel of Fig. 1(e). This abrupt change indicates the presence of a well-defined Fermi surface, with the position of this sudden change characterizing the Fermi momenta. These observations are robust for the strength of intermediate-to-infinite interactions $U/t$ and different system sizes $N$, suggesting that the electrons in the ferromagnetic phase for both intermediate $U/t$ and infinite $U/t$ limit are itinerant.
Kinetic Mechanism for Itinerant Ferromagnetism.— To understand the underlying mechanism for the itinerant ferromagnetism, we examine the distinct microscopic hopping processes encompassed in the hopping term $H_{t}$ of the intermediate-$U$ Hubbard model, as schematically illustrated in Figs. 2(a)-(c). Although all these hopping processes are allowed, whether ferromagnetism is induced specifically by a single process or several cooperative processes has yet to be conclusively established. We introduce the projection operator $\hat{Q}_{\mathbf{i},{\sigma}}\equiv n_{\mathbf{i},{\sigma}}$ and $\hat{P}_{\mathbf{i},{\sigma}}\equiv 1-\hat{Q}_{\mathbf{i},{\sigma}}$ to decompose the hopping term $H_{t}$:
${\small\begin{split}H_{t}=&-\sum_{\langle\mathbf{i}\mathbf{j}\rangle,\sigma}\bigl{(}t_{\mathrm{PQ}}\hat{P}_{\mathbf{i},\bar{\sigma}}c_{\mathbf{i},\sigma}^{\dagger}c_{\mathbf{j},\sigma}\hat{Q}_{\mathbf{j},\bar{\sigma}}+t_{\mathrm{QP}}\hat{Q}_{\mathbf{i},\bar{\sigma}}c_{\mathbf{i},\sigma}^{\dagger}c_{\mathbf{j},\sigma}\hat{P}_{\mathbf{j},\bar{\sigma}}\\ &+t_{\mathrm{PP}}\hat{P}_{\mathbf{i},\bar{\sigma}}c_{\mathbf{i},\sigma}^{\dagger}c_{\mathbf{j},\sigma}\hat{P}_{\mathbf{j},\bar{\sigma}}+t_{\mathrm{QQ}}\hat{Q}_{\mathbf{i},\bar{\sigma}}c_{\mathbf{i},\sigma}^{\dagger}c_{\mathbf{j},\sigma}\hat{Q}_{\mathbf{j},\bar{\sigma}}+\text{H.c.}\bigr{)}\\ \equiv&-\sum_{\langle\mathbf{i}\mathbf{j}\rangle,\sigma}\bigl{[}t_{\mathrm{PQ}}(H_{\mathrm{PQ}}+H_{\mathrm{QP}})+t_{\mathrm{PP}}H_{\mathrm{PP}}+t_{\mathrm{QQ}}H_{\mathrm{QQ}}\bigr{]}.\end{split}}$ (3)
Consequently, $H_{t}$ includes NN hopping between (i) a doubly occupied site and an empty site with hopping amplitude $t_{\mathrm{PQ}}$, or two singly occupied sites with hopping amplitude $t_{\mathrm{QP}}$ [see Fig. 2(a)]; (ii) a singly occupied site and an empty site with hopping amplitude $t_{\mathrm{PP}}$ [see Fig. 2(b)], i.e., singlon hopping; (iii) a doubly occupied and a singly occupied site with hopping amplitude $t_{\mathrm{QQ}}$ [see Fig. 2(c)], i.e., the doublon-singlon exchange. Given that case (i) describes Hermitian conjugate processes for an isotropic case, we consider $t_{\mathrm{PQ}}=t_{\mathrm{QP}}$. The hopping term in the Hubbard model (S1) corresponds to the isotropic limit $t_{\mathrm{PQ}}=t_{\mathrm{QP}}=t_{\mathrm{PP}}=t_{\mathrm{QQ}}=t$. This decomposition enables us to examine the individual contributions of each microscopic hopping to itinerant ferromagnetism.
As shown in Figs. 2(e-f) for $L_{y}=4,6$ DMRG cylinders, before entering the fully polarized region where $m_{s}^{2}(\bm{\Gamma})$ saturates, multiple hopping processes have nonzero expectation values, suggesting that all contribute to the ground-state kinetic energy. In contrast, when increasing $U/t$, only $\langle H_{\mathrm{QQ}}\rangle$ is nonzero in the fully polarized region. This observation strongly suggests that, among distinct hopping processes, the doublon-singlon exchange uniquely drives the fully polarized ferromagnetism. Then we focus on this hopping process and further analyze the role of lattice geometric frustration in resulting ferromagnetism.
The effective NN spin interactions $\mathbf{S}_{i}\cdot\mathbf{S}_{j}$ stems from two kinds of virtual hopping processes at finite electron doping: (i) the superexchange $t^{2}/U$ favoring antiferromagnetic order, (ii) the doublon-assisted ring-exchange $t^{3}/V^{2}$ [see Fig. 2(d)] favoring either antiferromagnetic or ferromagnetic order depending on the sign of hopping $t$ Thouless (1965); Tasaki (1998a), where $V$ refers to the on-site potential pinning the doublon. Therefore, the effective spin interaction is $\sim\lim_{V\to 0}(t^{2}/U+{t^{3}}/{V^{2}})\mathbf{S}_{i}\cdot\mathbf{S}_{j}$. This formulation reveals a fundamental difference between hole and electron doping in triangular lattices. For hole-doping ($t>0$), both hopping processes favor antiferromagnetism; while for electron-doping ($t<0$), the ring-exchange virtual process promotes ferromagnetism and finally dominates over the superexchange coupling in the large and infinite $U$ limit. Notably, square lattices lack such particle-hole asymmetry due to the positive ring-exchange coupling $\sim t^{4}$. This distinction highlights the important properties of geometrically frustrated lattices like the triangular lattice, where kinetic energy and geometric frustration work synergistically to foster itinerant ferromagnetism.
Fig. 3: Impact of the microscopic hopping processes on itinerant ferromagnetism in the Hubbard model. (a)-(f) Static spin structure factor $S(\mathbf{q})$ ((a)-(c)) and the momentum distribution $n(\mathbf{k})$ ((d)-(f)) obtained by independently turning off each hopping process: (a),(d) $t_{\mathrm{PQ}}=0$, (b),(e), $t_{\mathrm{PP}}=0$, (c),(f), $t_{\mathrm{QQ}}=0$. Other hopping terms in (3) are set to $t$. The white hexagon indicates the first Brillouin zone; interpolation is applied to contours.
To further verify the unique role of the doublon-singlon exchange in driving ferromagnetism, we independently deactivate each hopping amplitude while keeping the others unchanged, and compare the resulting ground states with the standard Hubbard model (S1). Therefore, the specific hopping process whose deactivation eliminates itinerant ferromagnetism is crucial for ferromagnetism. Figures 3 reveal that only switching off $t_{\mathrm{PQ}}$ or $t_{\mathrm{PP}}$ does not significantly affect the ferromagnetism, as the resulting $S(\mathbf{q})$ [see Figs. 3(a),(b)] and $n(\mathbf{k})$ [see Figs. 3(d),(e)] remain similar to those of the Hubbard model (S1) [see Figs. 1(c),(e)]. Conversely, switching off $t_{\mathrm{QQ}}$ eliminates ferromagnetic signatures [Figs. 3(c),(f)], since both the characteristic peak in $S(\mathbf{q})$ and well-defined Fermi surface in $n(\mathbf{k})$ disappear. These numerical results suggest the microscopic hopping process $H_{\mathrm{QQ}}$, as illustrated in Fig. 2(c), is pivotal in driving ferromagnetism at finite $U$. We further confirm this conclusion by examining a $t_{\mathrm{QQ}}-U$ model with only $t_{\mathrm{QQ}}=t$ and finite $U$ while $t_{\mathrm{PQ}}=t_{\mathrm{QP}}=t_{\mathrm{PP}}=0$, and the itinerant ferromagnetism is observed (details see SupplementarySM ).
Evolution of Magnetism with Doping.— Below, we explore magnetism with a broader doping range. We examine both the local and global spin correlations through the NN spin correlations and the static spin structure factors, respectively.
Fig. 4: Phase diagram. (a) Average NN spin correlations $\bar{S}(r=1)$ as a function of doping $\delta$ on width $L_{y}=4,6$ cylinders, resembling cold-atom quantum simulator findings Xu et al. (2023) for the same model. (b) Renormalized ferromagnetic order parameter $m_{s}^{2}(\mathbf{\Gamma})/m_{s,\mathrm{F}}^{2}(\mathbf{\Gamma})$ for the Hubbard model (S1) as a function of doping $\delta$ and interaction $U/t$. $m_{s,\mathrm{F}}^{2}(\mathbf{\Gamma})\equiv[(1-\delta)/2]^{2}$ is full polarization value of $m_{s}^{2}(\mathbf{\Gamma})$. Circles denote the data obtained numerically.
The average local spin correlations are defined as $\bar{S}(r)=\frac{1}{z}\sum_{\mathbf{j},|\mathbf{i}_{0}-\mathbf{j}|=r}\langle\mathbf{S}_{\mathbf{i}_{0}}\cdot\mathbf{S}_{\mathbf{j}}\rangle,$ with the summation over neighbors at a distance $r$ from a fixed site $\mathbf{i}_{0}$ that is strategically positioned at the lattice’s center to avoid a boundary effect. Here, we consider $r=1$ for NN spin correlations, and $z=6$ represents the coordination number of the triangular lattice. As shown in Fig. 4(a), the average NN correlations are antiferromagnetic across a broad range of hole doping ($\delta<0$) at intermediate coupling strength $U/t=10$, yet they are markedly suppressed by electron doping ($\delta>0$). In particular, we observe that the local spin correlations become ferromagnetic at intermediate electron doping. Notably, these local spin correlations at zero temperature align well with the recent cold-atom experimental findings (Fig. 3b of Ref. Xu et al. (2023)) at finite temperature for the similar intermediate $U/t$. This observation suggests that the local magnetic properties of the ground state survive in a finite range of temperatures.
We further examine the static spin structure factors to establish the magnetic phase diagram as a function of interaction $U/t$ and electron doping $\delta>0$. As illustrated in Fig.4(b) for $L_{y}=6$ cylinders, the renormalized squared order parameter $m_{s}^{2}(\bm{\Gamma})/m_{s,F}^{2}(\bm{\Gamma})$ shows ferromagnetism in a finite range of dopings $\delta\approx 1/2$ and interactions $U\gtrsim 10$, which smoothly extends to the infinite interaction limit. These findings are consistent with $L_{y}=4$ cylinders SM . Together with the numerical evidence presented in the previous section, this demonstrates that the ferromagnetism at both finite and infinite $U$ around $\delta\approx 1/2$ is driven by the doublon-singlon exchange.
Moreover, for finite $U$ and dopings both below and above $\delta\approx 1/2$, the incommensurate spin density wave (iSDW) emerges SM for both $L_{y}=4$ and $L_{y}=6$, which separates the ferromagnetic phase at $n=3/2$ from the phases near the full-filling and half-filling limits. We remark that, unlike the consistent observations at finite coupling strength, at $U=\infty$, we find a smooth connection between the ferromagnetic phase at $\delta\gtrsim 0$ and $\delta\approx 1/2$ when $L_{y}=6$, but these two phases are separated by an iSDW phase when $L_{y}=4$ SM . In addition, for larger doping $15/8\lesssim\delta<$ 1, both $L_{y}=4$ and $L_{y}=6$ systems exhibit another iSDW phase [see supplementarySM ], separating the ferromagnetic regions observed at $\delta\approx 1/2$ from ferromagnetism at $\delta\to 1$, thus suggesting distinct underlying mechanisms in these two doping concentrations.
Summary and outlook.— In this work, we explore the ferromagnetism mechanism of the finite-doped triangular-lattice Hubbard model at intermediate on-site Coulomb interaction. By diagnosing each microscopic hopping process separately, we reveal that the geometrically frustrated doublon-singlon exchange among other hopping processes in triangular lattices solely contributes to the fully polarized ferromagnetism. We further establish the magnetic phase diagram for the electron-doped triangular Hubbard model and compare the local spin correlations with the latest cold-atom experiments. Our findings may stimulate future studies of the triangular-lattice ferromagnetism beyond the cold-atom systems, particularly in condensed matter physics. For instance, the TMD moiré materials offer platforms with additional tuning parameters to study the stability of ferromagnetism against long-range Coulomb interactions in TMD hetero-bilayers Jin et al. (2021); Xu et al. (2020); Senthil et al. (2004a, b); Senthil (2023) and complex hopping in TMD homo-bilayers Wang et al. (2020). It may also inspire future finite-temperature numerical works to understand the relationship between rising temperature and increasing coupling strength, as previously focused on the half-filling case Wietek et al. (2021); Zhang et al. (2022).
Acknowledgements.
We particularly thank helpful discussions with Ashvin Vishwanath. This work is supported by the National Natural Science Foundation of China (Grant No.12074375), the Fundamental Research Funds for the Central Universities and the Strategic Priority Research Program of CAS (Grant No.XDB33000000).
Supplementary Materials for
“Ferromagnetism Mechanism in a Geometrically Frustrated Triangular Lattice”
I DMRG Details and DMRG Convergence
Fig. S1: Line-cut plot of the static spin structure $S(\mathbf{q})$ along the momentum path passing through the $\Gamma$ point (denoted as the white dashed line in Fig. 1(c) in the main text) with respect to various bond dimensions $D$.
In this section, we provide more numerical details in our calculations, as well as DMRG convergence with bond dimension. We study the ground-state properties of the triangular-lattice Hubbard model using DMRG. The Hamiltonian reads
$H=-t\sum_{\langle\mathbf{i}\mathbf{j}\rangle,\sigma}\left(c_{\mathbf{i},\sigma}^{\dagger}c_{\mathbf{j},\sigma}+\text{H.c.}\right)+U\sum_{\mathbf{i}}n_{\mathbf{i},\uparrow}n_{\mathbf{i},\downarrow}.$ (S1)
[…]
We employ DMRG to simulate the Hamiltonian in Eq. (S1) on triangular lattice cylinders with primitive vectors $\bm{e}_{x}=(1,0)$, $\bm{e}_{y}=(1/2,\sqrt{3}/2)$, wrapping on cylinders with a lattice spacing of unity. System size is $N=L_{x}\times L_{y}$, where $L_{x}$ ($L_{y}$) represents the length of the cylinder (circumference). Here we mainly study cylinders with width $L_{y}=4$, $6$ and set the energy unit as $t$. In particular, we extend our computation from finite-$U$ to the infinite-$U$ limit using the effective $t$-$J$ model with vanishing superexchange interaction $J=4t^{2}/U$. Considering the varying convergence rates at different parameters, we set the DMRG bond dimension $D$ up to $D\approx 36,000$.
We provide specific examples to illustrate the necessity of large bond dimensions for accurately identifying the ground-state properties of both the ferromagnetic phase and incommensurate SDW phases. Additionally, we have performed DMRG calculations with different initial states to ensure the convergence of DMRG.
For incommensurate SDW region, at relatively small bond dimensions, the peak of the static spin structure $S(\mathbf{q})$ is located at the momentum $\mathbf{\Gamma}$ [see Fig. S1(a,c,d)], which is a signature for ferromagnetism. However, with the increase of bond dimension, the energy becomes lower and the peak splits around the momentum $\mathbf{\Gamma}$. These observations indicate large bond dimensions are required to obtain converged results for distinguishing the ferromagnetism and iSDW with close energies.
On the other side, for the ferromagnetism region, a large bond dimension is crucial for determining whether the ground state is partially polarized or fully polarized. As presented in Fig. S1(b), the peak of $S(\mathbf{q})$ gets higher with increasing bond dimension, until it saturates. Notably, DMRG convergence becomes more challenging with decreasing interaction strength $U/t$ and increasing system size $N$. In Fig. S1(f,e), we also present the good convergence of the DMRG simulations in the iSDW region at $U=\infty$ that separates the ferromagnetic phases at $\delta\approx 1/16$ and $\delta\approx 1/2$ [see Fig. S6(b)].
II Details on unrestricted Hartree-Fock method
We consider the Hubbard model on the triangular lattice with the Hamiltonian
$H=-\sum_{\langle\mathbf{ij}\rangle,\sigma}(t+\mu\delta_{ij})c_{\mathbf{i},\sigma}^{\dagger}c_{\mathbf{j},\sigma}+U\sum_{i}n_{\mathbf{i}\uparrow}n_{\mathbf{i}\downarrow}$ (S2)
Here $t$ is the hopping integral and $U$ is the on-site Hubbard interaction strength. Our focus is on a scenario of high electron doping $\delta=0.5$ and applying the unrestricted Hartree-Fock (UHF) approximation. In the UHF approximation, the mean-field order parameters are introduced for both the on-site densities $\langle n_{\mathbf{i},\sigma}\rangle$ and the spin-flips $\langle S_{\mathbf{i}}^{-}\rangle$ and $\langle S_{\mathbf{i}}^{+}\rangle$. Then, the Hartree-Fock factorizes the Hamiltonian to
$H_{\mathrm{HF}}=$ $-\sum_{\langle\mathbf{ij}\rangle}(t+\mu\delta_{\mathbf{ij}})c_{\mathbf{i},\sigma}^{\dagger}c_{\mathbf{j},\sigma}+\sum_{i}U[\langle n_{\mathbf{i},\uparrow}\rangle n_{\mathbf{i},\downarrow}+\langle n_{\mathbf{i},\downarrow}\rangle n_{\mathbf{i},\uparrow}-\langle S_{\mathbf{i}}^{+}\rangle S_{\mathbf{i}}^{-}-\langle S_{\mathbf{i}}^{-}\rangle S_{\mathbf{i}}^{+}]+\sum_{\mathbf{i}}U\left[-\langle n_{\mathbf{i},\uparrow}\rangle\langle n_{\mathbf{i},\downarrow}\rangle+\langle S_{\mathbf{i}}^{-}\rangle\langle S_{\mathbf{i}}^{+}\rangle\right],$
Here the mean-field parameters is an expectation value under the ground state $|\mathrm{HF}\rangle$ of the $H_{\mathrm{HF}}$ mean-field Hamiltonian,
$\langle n_{\mathbf{i},\sigma}\rangle\equiv\langle\mathrm{HF}|n_{\mathbf{i},\sigma}|\mathrm{HF}\rangle,\quad\langle S_{\mathbf{i}}^{\pm}\rangle\equiv\langle\mathrm{HF}|S_{\mathbf{i}}^{\pm}|\mathrm{HF}\rangle,$ (S3)
and the on-site densities are subject to the particle number constraint $\sum_{\mathbf{i}}[\langle n_{\mathbf{i},\uparrow}\rangle+\langle n_{\mathbf{i},\downarrow}\rangle]=N(1+\delta)$, with $N$ being the number of the lattice sites. We have total $4N$ variational parameters, which can be iteratively solved by the self-consistent equations
$\langle n_{\mathbf{i},\sigma}\rangle$ $=\sum_{n}f_{FD}(E_{n})\langle\psi_{n}|n_{\mathbf{i},\sigma}|\psi_{n}\rangle$ (S4)
$\langle S_{\mathbf{i}}^{\pm}\rangle$ $=\sum_{n}f_{FD}(E_{n})\langle\psi_{n}|S_{\mathbf{i}}^{\pm}|\psi_{n}\rangle$ (S5)
where $|\psi_{n}\rangle$ is an eigenstate of $H_{\mathrm{HF}}$ with $H_{\mathrm{HF}}|\psi_{n}\rangle=E_{n}|\psi_{n}\rangle$ and $f_{FD}(E_{n})$ is the Fermi-Dirac distribution.
Fig. S2: The static spin structure factor $S(\mathbf{q})$ as a function of $U/t$ at electron doping $\delta=1/2$. (a)-(c) The contour plot of $S(\mathbf{q})$ for the $N=24\times 6$ system at $U/t=4$ (a), $U/t=10$ (b), and $U/t=24$ (c). The black dots represent the accessible momenta in the Brillouin zone (denoted as a white hexagon), and interpolation has been applied in the contour plot. (d) Schematic of the evolution of the $\mathbf{q}_{0}$ with increasing $U/t$, indicated by the transition from lighter to darker colors following the black dashed arrows. Here, $\mathbf{q}_{0}$ denotes the momenta where $S(\mathbf{q})$ reaches its maximum, within the Brillouin zone represented by a hexagon. (e) Ferromagnetic squared order parameter $m_{s}^{2}(\mathbf{\Gamma})$ as a function of $U/t$. For larger $U/t$, the peak of $S(\mathbf{q})$ stabilizes at momentum $\mathbf{\Gamma}$ with a saturation value independent of $N$, indicating the fully polarized ferromagnetism. (f) Line-cut plot of $S(\mathbf{q})$ along the momentum path passing through the $\mathbf{\Gamma}$ point, depicted as a white dashed line in (c). (g) Polarization $P\equiv 2Nm_{s}(\mathbf{\Gamma})$ from the unrestricted Hartree-Fock method on $N=48\times 48$ and $N=24\times 24$ systems with periodic boundary conditions.
For a stable convergence, we utilize the direct inversion in the iterative subspace (DIIS) method. In this approach, we can express the Hartree-Fock Hamiltonian in matrix form as follows:
$H_{\mathrm{HF}}=C^{\dagger}FC+\sum_{\mathbf{i}}U\left[-\langle n_{\mathbf{i},\uparrow}\rangle\langle n_{\mathbf{i},\downarrow}\rangle+\langle S_{\mathbf{i}}^{-}\rangle\langle S_{\mathbf{i}}^{+}\rangle\right]$ (S6)
where $C$ is a vector with $C_{\mathbf{i}\sigma}=c_{\mathbf{i},\sigma}$ and $F$ is the Fock matrix. For the magnetic order, the single-particle correlation matrix $\Phi$ is defined as
$\Phi_{\mathbf{ij},\alpha\beta}=\langle c_{\mathbf{i},\alpha}^{\dagger}c_{\mathbf{j},\beta}\rangle.$ (S7)
The DIIS method relies on the property that a convergence condition is met when the single-particle correlation matrix commutes with the Fock matrix
$F\Phi-\Phi F=0.$ (S8)
This condition allows us to define an error vector for the $n$-th iteration as
$\mathbf{e}_{n}=F_{n}\Phi_{n}-\Phi_{n}F_{n},$ (S9)
where $\Phi_{n}$ is obtained from the diagonalization of $F_{n}$. The iteration process involves updating the Fock matrix as
$F_{k}=\sum_{n=1}^{k-1}d_{n}F_{n}.$ (S10)
Here, $d_{n}$ are the DIIS coefficients obtained through a least-squares constrained minimization of the error vectors. The aim of the optimization is to minimize the error given by
$\mathrm{Error}$ $=|\sum_{n}d_{n}\mathbf{e}_{n}|^{2}$ (S11)
with the constraint $\sum_{n}d_{n}=1$. In our calculations, we set the convergent condition of the energy difference between two subsequent iteration processes as $10^{-8}t$.
III Itinerant Ferromagnetism at Electron Doping $\delta=1/2$
Fig. S3: (a-d) The contour plot of the static spin structure factor $S(\mathbf{q})$ across a range of $U/t$ with $N=8\times 4$ at $\delta=1/2$ for the Hubbard model.
Fig. S4: (a-d) The contour plot of the static spin structure factor $S(\mathbf{q})$ across a range of $U/t$ with $N=12\times 4$ at $\delta=1/2$ for the Hubbard model.
In the main text, we have presented key findings at electron doping $\delta=1/2$. In this section, we provide additional numerical details and numerical data. For the convenience of readers, we also show the data presented in the main text to gain a complete picture of the evolution of magnetism when tuning the coupling strength $U/t$ from finite to infinite $U/t$, as depicted in Fig. S2. These analyses reveal the emergence of itinerant ferromagnetism with increasing Coulomb interaction and enable a comprehensive understanding of the underlying mechanism under general interaction strength $U/t$.
The magnetic order is signaled by peaks at certain wave vectors $\mathbf{q}=\mathbf{q}_{0}$ in the static spin structure factor $S(\mathbf{q})$, defined by $S(\mathbf{q})\equiv\sum_{\mathbf{i},\mathbf{j}}\langle{\mathbf{S}_{\mathbf{i}}\cdot\mathbf{S}_{\mathbf{j}}}\rangle e^{i\mathbf{q}\cdot(\mathbf{i}-\mathbf{j})}/N$. We first focus on the evolution of the momenta $\mathbf{q}_{0}$ as a function of interaction strength $U/t$. For the noninteracting case ($U/t=0$) in a triangular lattice, electron doping $\delta=1/2$ corresponds to the Van-Hove filling at which a perfect nesting of the Fermi surface exists. In the weak-coupling limit $U/t\to 0$, it is expected that nesting leads to peaks of $S(\mathbf{q})$ locating at the $\mathbf{M}$ points in the Brillouin zone. Previous weak-coupling theory proposes that the candidate ground state exhibits noncoplanar chiral magnetic order. When the weak coupling assumption is no longer applicable at intermediate to strong interaction strengths $U/t$, using DMRG we numerically observe that other magnetic ordering vectors distinct from the nesting vectors emerge, as shown in Figs. S2(a), (b) for $L_{y}=6$. Interestingly, the locations of these peaks in $S(\mathbf{q})$ shift towards the $\mathbf{\Gamma}$ point as $U/t$ increases. Upon surpassing a critical strength $U_{c}/t$ toward the limit $U/t=\infty$, the peak of $S(\mathbf{q})$ becomes fixed at $\mathbf{\Gamma}$ point, as illustrated in Fig. S2(c), (e), indicating the establishment of a stabilized ferromagnetic phase in the strong coupling regime. In Fig. S2(d), we schematically plot such evolution of the peaks in $S(\mathbf{q})$ with increasing $U/t$, as visualized by the dashed arrow from momenta $\mathbf{M}$ to $\mathbf{\Gamma}$. Notably, after the peak of $S(\mathbf{q})$ shifts to the $\mathbf{\Gamma}$ point, its height gets larger in a finite range of increasing $U/t$ [see Fig. S2(f)], until reaching its maximum value and forming a plateau, as illustrated by Fig. S2(e). This signals the evolution from a partially-polarized to a fully-polarized phase. In particular, the saturated value matches the ferromagnetic squared order parameter $m_{s}^{2}(\mathbf{\Gamma})\equiv S(\mathbf{\Gamma})/N\sim\delta_{\mathbf{q},\mathbf{\Gamma}}/16$ for a fully polarized state at electron doping $\delta=1/2$.
We present additional supporting data for the evolution of the momentum $\mathbf{q}_{0}$ of the peak in $S(\mathbf{q})$ as a function of interaction strength $U/t$ in Fig. S3 for $N=8\times 4$ and Fig. S4 for $N=12\times 4$. These results for $L_{y}=4$ systems are consistent with the case of $L_{y}=6$ in Fig. S2(a)-(c). The peak locations in $S(\mathbf{q})$ progressively shift toward the $\mathbf{\Gamma}$ point as $U/t$ increases, as presented in Fig.S3(a-c) and Fig.S4(a,b). This observation supports the schematic plot in Fig. S2(d). After the peak of $S(\mathbf{q})$ shifts to the $\mathbf{\Gamma}$ point, its height increases within a specific range of $U/t$ [see, Fig. S4(b-d)], until reaching its maximum value, as illustrated by Fig. S3(c,d) and Fig. S4(c,d), indicating a transition from a partially-polarized to a fully-polarized ferromagnetic phase.
The above numerical observations reveal the presence of ferromagnetism with increasing Coulomb interaction. To confirm its robustness in the thermodynamic limit, we further employ the unrestricted Hartree-Fock mean-field approximation (see details in Supplementary Materials Section II). As shown in Fig. S2(g), the polarization $P\equiv 2Nm_{s}(\mathbf{\Gamma})$ increases gradually as $U/t$ becomes larger, and finally a full polarization state occurs at $U\simeq 4.5t$ for both $N=48\times 48$ and $N=24\times 24$ systems under periodic boundary conditions on both directions. While disparities in polarization $P$ between $N=48\times 48$ and $N=24\times 24$ lattice sizes are notable for $U\lesssim 4.5t$, beyond this threshold ($U\gtrsim 4.5t$) in the fully polarized ferromagnetic state, the polarization $P$ shows minimal sensitivity to the size of the system. This echos the observations in Fig. S2(e), suggesting the robust fully-polarized ferromagnetism regardless of system sizes in the strong coupling regime.
Fig. S5: Itinerant ferromagnetism signature for the $t_{\mathrm{QQ}}-U$ model. (a) The static spin structure factor $S(\mathbf{q})$. (b) The electron momentum distribution $n(\mathbf{k})$. The black dots represent the accessible momenta in the Brillouin zone (white lines). Interpolation has been applied in the contour plot. Here, $N=8\times 4$, $U/t=10$, $\delta=1/2$.
Fig. S6: (a) The average spin correlations $\bar{S}(r=1)$, as defined in the main text and indicated by yellow triangles ($L_{y}=4$) and green diamonds ($L_{y}=6$), are calculated across the nearest-neighbor bonds for the standard Hubbard model. These results are compared with those (blue circles) obtained when $t_{\mathrm{QQ}}=t$ in Eq. (3) in the main text, with all other hopping processes absent. The dashed black line signifies the fully polarized case. The largest system sizes we presented here are $N=16\times 6$ and $N=18\times 4$ for 6- and 4-leg cylinders, respectively. (b) The renormalized ferromagnetic order parameter $m_{s}^{2}(\mathbf{\Gamma})/m_{s,\mathrm{F}}^{2}(\mathbf{\Gamma})$ for the Hubbard model as a function of doping $\delta$ and interaction $U/t$. Here, $m_{s,\mathrm{F}}^{2}(\mathbf{\Gamma})\equiv[(1-\delta)/2]^{2}$ corresponds to a full polarization value of $m_{s}^{2}(\mathbf{\Gamma})$. Circles denote the data obtained numerically.
Fig. S7: Contour plots of the static spin structure factor $S(\mathbf{q})$ for the Hubbard model, illustrating two distinct iSDW phases. Panels (a) and (c) show phases at smaller doping, while panels (b) and (d) depict phases at larger doping.
IV Itinerant Ferromagnetism of $t_{\mathrm{QQ}}-U$ model
The $t_{\mathrm{QQ}}-U$ model is given by $H=H_{t_{\mathrm{QQ}}}+H_{\mathrm{int}}$, where the interaction term is the Hubbard repulsion $H_{\mathrm{int}}=U\sum_{\mathbf{i}}n_{\mathbf{i}\uparrow}n_{\mathbf{i}\downarrow}$, and the projected hopping term is given by
$H_{t_{\mathrm{QQ}}}=-\sum_{\langle\mathbf{i}\mathbf{j}\rangle,\sigma}t_{\mathrm{QQ}}\hat{Q}_{\mathbf{i},\bar{\sigma}}c_{\mathbf{i},\sigma}^{\dagger}c_{\mathbf{j},\sigma}\hat{Q}_{\mathbf{j},\bar{\sigma}}.$ (S12)
Here, the projection operator is $\hat{Q}_{\mathbf{i},{\sigma}}\equiv n_{\mathbf{i},{\sigma}}$ and Eq. (S12) describes the NN hoppings between a doubly occupied site and a singly occupied site with hopping amplitude $t_{\mathrm{QQ}}$ [see Fig. 2(c) in the main text], i.e., the doublon-singlon exchange. In Fig. S5, we examine both the static spin structure factor $S(\mathbf{q})$ and electron momentum distribution $n(\mathbf{k})$ at electron doping $\delta=1/2$ for $U/t=10$, where the pronounced peak at $\Gamma$ and the well-defined Fermi surface are consistent with the itinerant ferromagnetism observed in the Hubbard model at the same doping and the same coupling strength $U/t=10$.
V Evolution of Magnetism with Doping for $L_{y}=4$
In this section, we present the average spin correlations $\bar{S}(r=1)$ and the magnetic phase diagram as a function of $U/t$ and electron doping $\delta$ specifically for $L_{y}=4$. We also provide further data for $L_{y}=6$ to enhance the comparison.
For the fully polarized phase of the Hubbard model with $U/t=10$, $\delta\approx 1/2$, $L_{y}=4$, the average spin correlations $\bar{S}(r=1)$ [yellow triangles in Fig. S6(a)] align well with those [blue circles] of the $t_{\mathrm{QQ}}-U$ model Eq. (S12). This alignment underscores a deep connection between the ferromagnetic mechanisms inherent to both cases, and thus highlights the significant influence of doublon-singlon exchange in inducing ferromagnetism in the Hubbard model with finite $U/t$. In addition, for light electron doping around $\delta\approx 1/16$, the $\bar{S}(r=1)$ of the $t_{\mathrm{QQ}}-U$ model [blue circles in Fig. S6(a)] closely matches with the fully polarized value (dashed black line) $m_{s,\mathrm{F}}^{2}(\mathbf{\Gamma})\equiv[(1-\delta)/2]^{2}$, yet deviation becomes apparent around a doping range of $43/16\lesssim\delta\lesssim 1/$ 4, the same as the ferromagnetic feature of the Hubbard model at $U/t=\infty$ for $L_{y}=4$, as shown by the Fig. S6(b).
Figure S6(b) presents the magnetic phase diagram as a function of $U/t$ and electron doping $\delta$ for $L_{y}=4$, identifying two distinct ferromagnetic phases at $U=\infty$ around doping $\delta\approx 1/16$ and $\delta\approx 1/2$, separated by a doping range of $43/16\lesssim\delta\lesssim 1/$ 4. The pattern of $S(\mathbf{q})$ at such separation [see Fig. S7(a)] is similar to that with finite $U$ [see Fig. S7(c)], indicating that the iSDW at small doping extends from finite $U$ to $U=\infty$ smoothly for the case of $L_{y}=4$. This numerical observation suggests the subtle competition between the iSDW and FM at intermediate doping on narrower cylinders like $L_{y}=4$ and the final stabilization of FM on wider cylinders like $L_{y}=6$ [see Fig. 4(b) in the main text]. The notable separations of ferromagnetic phases around $\delta\approx 1/16$, $\delta\approx 1/2$ and $\delta\to 1$ for a finite system with $L_{y}=4$ hint at diverse underlying mechanisms, potentially differentiating the observed ferromagnetism around $\delta\approx 1/2$ from the Nagaoka ferromagnetism with $\delta\to 0$ and the Mielke’s flat-band ferromagnetism or Müller-Hartmann ferromagnetism with $\delta\to 1$ for a finite size system, even though such separations of the ferromagnetic phases are absent for $L_{y}=6$ [see Fig. 4(b) in the main text] at least for the largest length $L_{x}=16$ that we have simulated. On the other hand, as the effective interaction $U/t$ decreases from $\infty$ to $U/t=10$ for $L_{y}=4$, the ferromagnetic region around lighter ($\delta\approx 1/16$) and intermediate ($\delta\approx 1/2$) doping disappears and shrinks [Fig. S6(b)], respectively, yielding a ferromagnetic landscape closely aligned with the $U/t=10$ case as depicted in Fig. S6(a). Additionally, for larger doping $15/8\lesssim\delta<$ 1, a distinct pattern in $S(\mathbf{q})$ indicates the presence of another iSDW phase [see Fig. S7(b,d)].
We further present the $S(\mathbf{q})$ outside the ferromagnetism phase at small doping with finite $U$ and larger doping $15/8\lesssim\delta<$ 1 in Fig. S7. Our examination reveals two distinct features in $S(\mathbf{q})$, each indicative of a different iSDW phase.